# Gravitational Wave emission in Binary Neutron Star early post-merger within a dark environment

### v1

@@ Abstract
Using an effective Lagrangian model inspired by [1] we describe the early post-merger of a binary Neutron Star (BNS) merger event with a non-vanishing ambient fraction of dark matter. For this we mimic the dynamics of two oscillating Neutron Star (NS) masses in the gravitational potential well as they merge. We parametrize the ejected baryonic matter fraction in the coalescence event forming a surrounding debris disk containing a non-vanishing dark matter fraction. In order to analyze the dark contribution, we start from a dark-matter free modelization using Monte Carlo Markov Chain (MCMC) techniques to replicate the gravitational waveforms obtained from simulations in the CoRe database for three selected equations of state for NS matter (DD2, BHBlp, SFHo). Later, we explore the impact of a dark viscous fluid under a prescribed velocity dependent force and obtain the waveforms and spectral properties originating in the first few ms in the BNS post-merger. Finally we discuss our findings in light of the prospects of detectability of a non-vanishing fraction of dark matter arising in the waveforms and power spectral distribution in present or future experimental settings.
dark matter, gravitational waves, binary neutron star merger, effective lagrangian, spectral properties
@@ I Introduction
So far different particle candidates have been proposed to account for the DM. Ranging from weakly to strongly interacting species and for a mass range spanning many orders of magnitude, there are plenty of models in the literature. Just to cite a few of them, let us remind here the weakly interacting massive particle (WIMP) such as the supersymmetric neutralino, asymmetric dark matter (ADM), light bosons such as axions or axion-like particles and strangelets [6]. Although much theoretical and experimental efforts have been undertaken, its nature remains elusive, constituting an open problem. Nevertheless, there is now overwhelming evidence of its existence more than nine decades after the pioneering observations by Fritz Zwicky in 1933 [7], who showed that the velocities of galaxies in the Coma cluster greatly exceeded the expectations based solely on the sum of the individual luminous galaxy masses. Since then, wide searches involving different strategies trying to hunt these exotic particles have shown null results although the phase space is now more tightly constrained.
@@
Remarkably, the dawn of the era of multimessenger physics allows to gather information on different channels enriching the picture and more efficiently constraining the physical mechanisms governing its behaviour. In particular, gravitational waves (GWs) experiments are expected to provide a precise reconstruction of the properties of inspiraling and merging compact objects, including mass/radius/composition measurements as they would be sensitive to tiny deviations in the gravitational waveforms which may be induced by matter features in the violent events of the merging black holes (BHs) and NSs. These environmental effects could be the smoking-gun to signal the presence of DM, although this requires a careful identification of the ordinary baryon content [8].
@@
Since the LIGO first detection in 2015 of GWs emitted during the merger of a stellar-mass binary black hole (BBH) merger event [9], the possibility to add this new multimessenger opens a new window to the Universe. In addition, GWs provide a new channel to shed light on the nature of DM, also constituting a definite confirmation that BBHs exist and merge with a local rate [10] in the range between 17.9 and $44\hskip 2.0pt\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ at a fiducial redshift $(z=0.2)$, assuming the known sources are representative of the total population. Other compact binary coalescence (CBC) search targets include collisions of BHs and NSs or a binary NS merger(BNS), where in all of them GWs are produced in a transient emission that can be detected on current terrestrial facilities. Recently, detectors including LIGO-Virgo-Kagra (LVK) have released enough data to build catalogs currently listing nearly 100 detections [11]. Specifically the Gravitational-Wave Transient Catalog 3 (GWTC-3) [10] contains signals consistent with three classes of binary mergers. They infer the BNS merger rate to be between 10 and $1700\hskip 2.0pt\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ and the NS-BH merger rate to be between 7.8 and $140\hskip 2.0pt\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, assuming a constant rate density in the comoving frame and taking the union of $90\%$ credible intervals. Besides these efforts, other future 3rd-generation planned detectors such as Cosmic Explorer (CE) [12], Einstein Telescope [13], NEMO[14] or LISA[15] will be able to further detect ultralow and ultraweak GW spectral sensitivities of the order $h\sim 10^{-25}$ $\rm Hz^{-1/2}$ produced by massive cosmic events with redshifts $z\lesssim 10$ and total mass around that expected in BNS $M\lesssim 3M_{\odot}$.
[…]
@@
In this work we will use a Lagrangian effective model inspired by [1, 17] taylored to study the dynamics of the early post-merger phase in a BNS merger occurring in an environment polluted by viscous DM. Thus in our model the BNS environment posses a baryonic mass fraction $f_{b}$ and DM fraction $f_{\chi}$ during the GW emission. In this way we are able to compute the set of polarization strengths $h_{+,\times}(t)$ of the GW signal as a function of time $t$ and subsequent observables derived from them. We note that, at present, there are additional tools such as Bilby [18], a Bayesian inference library for gravitational-wave astronomy, allowing to infer the source astrophysical properties from existing experimental data. This tool does not incorporate, however, particular realizations of exotic models including DM like the one we present in this work.
@@
In our setting the presence of DM in the system is due to the actual BNS merger having place in a DM polluted environment. It may be due to stripped DM from the inside of the NS or from an overdensity or halo where typically DM density can be much larger with respect to the solar neighbourhood value $\rho_{\chi}\simeq 0.4$ $\rm GeV/cm^{3}$. It has been recently claimed [19] that even overdensities of a few $\sim 10^{4}$ $\rm GeV/cm^{3}$ at 1 parsec from the center of M4 would not yield definite constraints on DM from kinetically heating of NSs or white dwarfs.
@@
The detection of DM using GW along with different strategies has already been discussed, see [8]. One of the key ideas is based on the fact that it constitutes a fingerprint of the formation and evolution of first cosmic structures. Popular models include cold DM (CDM) on sub-galactic scales. However, there still remain incompatibilities between the CDM description and the observed data, as it predicts more structure on small scales than what we actually observe. Other models such as warm or fuzzy DM, as shown by simulations, lead to the suppression of small-scale structures and this, in turn, may lead to a significant reduction in the rate of BBH mergers observed, helping to further constrain among existing DM models. On the small scales, collisionless DM fluid faces major issues such as the core cusp problem, missing satellites problem, too big to fail problem. For more detail on the small scale issues, see [20]. This shows that future GW observations will provide a new probe of physics beyond the $\Lambda$CDM model [21].
@@
The possible traces that many of the DM candidates could leave on the patterns of GW polarizations has been a particularly effervescent topic in the last decade [22, 23, 24, 25, 26]. The existence of viscous and self-interacting DM (SIDM) is another key point. At the small scale, the SIDM behaves like a collisional dark matter but on the large scale it behaves like the collisionless DM, thus fitting expectations. In [27] it is pointed out that the collisional nature of DM can lead to viscosity and using the kinetic theory and viscous coefficients of the SIDM can be obtained. Therefore the ratio of the SIDM cross section to its mass $(\sigma/m_{\chi})$, for the present observed cosmic acceleration, can be reconciled with the constraints obtained from astrophysical observations. Note that a GW will also experience the damping effect when it propagates in a fluid with non-zero viscosity so that the constraints of the luminosity distances from the observed GW events by experiments can be directly used to set constraints on self-interactions [28, 29].
@@
The structure of this contribution is as follows. In section II the Lagrangian model is introduced along with the relevant set of parameters necessary to describe the GW polarizations and subsequent quantities obtained from them. In our description, inspired by the mechanics of a two-body oscillator, we include typical BNS merger associated quantities, such as the mass ratio and the chirp mass, aimed to effectively describe the GW emission. We discuss how the presence of a hybrid viscous medium allows mimicking selected GW catalog waveforms during the $\sim 3$ ms timescale after the merger. Thus we will compare signals from recent numerical relativity simulations from the Computational Relativity (CoRe) database [30] involving several equations of state (EoS). In section II.1, the different polarization modes in the GW from the selected dataset are fit, first to a DM-free model, in order to assess the efficiency of the Monte Carlo Markov Chain (MCMC) based fitting procedure allowing for a parameter estimation and, later on, allowing for a non-vanishing viscous DM fraction. We provide a physical insight of the effective parameters used.
@@
In section III we calculate the power spectral density (PSD) in order to characterize the frequency domain of the expected GW features and the sensitivity of current and future detectors to novel dark components, based on the relative effective signal-to-noise ratio for different DM polluted environments. Finally in section IV we conclude.
@@ II Model for the emission of Gravitational waves
First we will introduce the Lagrangian model that we will use in order to calculate the GW emission in the post-merger phase of the BNS event. The system we will scrutinize are the first few ms right after the NSs merge. Note that in this phase, strong tidal forces have already disrupted the otherwise nearly spherical (rotating) stellar configuration in asymptotically early times before the merger. In our scenario we will assume that a dark polluted environment is present due to a non-vanishing fraction of DM coexisting with the expected baryonic matter mass fraction expelled in the collision. This dark fraction origin may well come either from a liberated dark population loosely bound in superficial stellar layers or beyond baryonic radii, such as in halo configurations, see [4] or from an existing local DM overdensity if the BNS merger takes place inside a DM cluster.
As computational simulations show [31], for the few ms duration the BNS early post-merger dynamics can be mimicked by that of a damped oscillator (DO). This notion arises naturally as the GW transient emission takes place in a baryon environment with associated dissipative oscillating dynamics from ordinary (nuclear) matter.
@@
In this scenario, we study how a fluid composed of massive $\chi$ particles may leave a fingerprint in the transient emission of the BNS merger event we model, using a minimally coupled DM model. In this respect DM should be thermalized with the NS stellar matter so that it can (at least partially) follow the dynamics of the merger. This condition is fulfilled for DM models yielding constant scattering spin-independent $\chi$-nucleon cross-sections for $m_{\chi}\gtrsim 1$ GeV$/\rm c^{2}$, $\sigma_{\chi,N}\gtrsim 10^{-45}$ $\rm cm^{2}$, see [32]. Further couplings at the strong (hadronic) force limit are theoretically possible yielding a more ultradense environment [33] driving the final object to a BH collapse. However, we will restrict the DM sector we consider to yet another fraction of matter component, essentially inert with respect to changes in ordinary baryonic matter species. Thus in the post-merger the stellar configuration corresponds to that of two stellar nucleus surrounded by a disk (assuming some composition of previous spin from the merging stars) with baryonic and DM content. The actual proportions of the total mass in the disk and nuclei are to be determined dinamically, see below.
II.1 Binary Neutron Star merger as a damped oscillator
As mentioned, we describe this early post-merger as a DO immersed in ejecta material. For simplicity we will use a viscous (dissipative) force $F_{b,d}\sim v_{r}^{n}$ as a power of radial velocity $v_{r}=\dot{r}$. We will consider that the DO dynamics are affected by this non-ideal behavior in two ways, first from a nuclear (baryon) fluid with associated quadratic viscous resistance $F_{b,d}\sim-\alpha_{2}\dot{r}^{2}$ while a less dense, diffuse DM fraction contributes instead $F_{\chi,d}\sim-\beta_{n}\dot{r}^{n}$. $\alpha_{2},\beta_{n}\,(n=1,2)$ are effective parameters characterizing both types of matter that will be discussed later.
@@ II.1 Binary Neutron Star merger as a damped oscillator
Using MCMC we will fit the oscillation modes as a function of time, obtained from simulations such as CoRe database [30] for a set of three different EoS, namely BHBlp [34], SFHo [35], and DD2 [36]. All of them can describe NS matter with a non-zero viscosity and finite temperature. As described in [37, 31], immediately after the merger, peak temperatures $T\sim 80$ MeV are reached. We note the dark fluid we introduce will nevertheless be a small mass fraction with a proper viscosity that may be lower or comparable to that of baryons. As discussed, the ultimate source of DM viscosity is an assimilated non-zero self-interaction or $\chi$-nucleon cross-section that although weak in strength may be, however, similar to photon-photon scattering in electromagnetic plasmas for energies or hundredths of MeV.
@@
More specifically dark and baryonic matter will be expelled from the NS outer layers, less gravitationally bound, due to the tidal forces. In our setting we will consider this fraction of ejected matter forms a disk around the NS nuclei rotating with the same angular velocity. Note that this picture is what current numerical relativity simulations yield [38] and the scenario for kilonova emission, detectable with Earth based telescopes such as MAAT [39], in the neutron-rich environment formed after the merger.
@@
We implement this scenario by means of two merging NSs with gravitational masses $m_{1}$ and $m_{2}$ in a rotating disk with a fraction of DM, $f_{\chi}$, and an ordinary i.e. baryonic fraction, $f_{b}$, whose origin is the debris of the tidal disruption of the two compact stars. In the mechanical analog, both spheres are connected by a spring characterized by constant $k$. The spheres can oscillate along the radial coordinate and rotate on the two-dimensional orbiting plane. For the sake of simplicity, we will consider an observer perpendicular to the plane where the orbital motion happens. As the damping oscillation proceeds, there will be a dynamical variation of the moment of inertia of the system and, from the conservation of angular momentum, this will be resulting in the generation of GW with a specific spectrum.
@@
This model is inspired by previous works [1, 17] describing, however, a different case i.e. a four-mass oscillator where two DM and two NS matter cores oscillate jointly with two different springs with constants $k_{1},k_{2}\neq k_{1}$. Our model is different as it considers the fact that the oscillations happen in a dark environment and aims at exploring the spectral features observed in the early post-merger signal. As we will describe, the frequencies and amplitudes of these oscillations will be affected by the presence of DM and this, in turn, determines the spectral properties of the GW signal.
Despite its simplicity, this effective model can reproduce the complex waveforms emitted in the early BNS merger and provide insights into the physical interpretation of the various peaks in the spectrum. It provides a convenient tool to improve post-merger signal description and modeling, potentially enabling the development of comprehensive analytical models for template construction including a DM component.
@@
In order to evaluate the sensitivity of our model to DM in the predicted GW signal strength we will keep restricted to a minimum the number of extra parameters describing the dark environment where the BNS happens. As we will explain, we include namely the DM mass fraction $f_{\chi}$ and one physical parameter, a dissipative coefficient, $\beta_{n}$, describing the radial velocity ($\dot{r}$) dependent viscous force $F_{\chi,d}\sim\beta_{n}v^{n}$, characterizing the presumable viscosity of the low-density DM fluid. Note that this effectively non-relativistic parametrization is based on the fact that there may be non-vanishing cross-sections among the two sectors and a possibly self-interacting term in the dark sector. As we will show, the non-relativistic nature of DM dynamics can be included in our minimal scenario in a meaningful way, yielding characteristic features capable of constraining some DM models.
@@
As mentioned above, the Lagrangian we consider is able to reproduce not only the GW polarizations but also many spectral properties of the BNS early post-merger signal. Thus the BNS merger event can be thought of in this scenario as two phases. In a first phase of the merger the outer NS layers are stripped off due to the tidal fields, carrying away the baryonic mass $m_{b}=f_{\text{ejec}}(m_{1}+m_{2})\equiv f_{\rm ejec}M_{T}$ from the progenitor stars ($m_{1},m_{2}$) rapidly forming a disk-like structure with a radius $R_{d}$, enclosing the two NS inner nuclei of masses $(1-f_{\text{ejec}})m_{1}$ and $(1-f_{\text{ejec}})m_{2}$, respectively bound by an elastic force of constant $k$. We use equivalently $f_{\text{ejec}}\equiv f_{b}$. The minimum elongation happens at a characteristic radial distance, $l$, from the coordinate origin. The original motivation of this oscillating behaviour is guided by numerical relativity simulations showing this effect in a gravitational potential well created during the merger phase [40]. Thus we will be assuming that the existence of this elastic force is an approximation that captures the behavior of these two NS nuclei and their motion in a quasi-elliptical orbit.
Due to the abrupt fall of the tidal field after the disruption of the outer layers and the conservation of angular momentum, these nuclei keep rotating inside this pre-collapsed object with the same angular velocity $\omega\equiv\dot{\theta}$, which therefore has a metastable nature. Later on, a complete fusion is occurring inside (second phase) as soon as the rotational energy of these nuclei decays due to dissipation through interactions with the surrounding matter in the radial direction.
@@
For a system of mass ratio $q=m_{1}/m_{2}$, we can write the following Lagrangian as ideal starting point

### v2

@@ Abstract
Using an effective Lagrangian model inspired by Takami et al. [1] we qualitatively study the early post-merger of a nearly symmetric binary Neutron Star (BNS) merger event with a non-vanishing ambient fraction of dark matter. For this we first mimic the dynamics of two oscillating Neutron Star (NS) masses in the gravitational potential well as they merge. We parametrize the dynamics and ejecta properties in the coalescence event allowing the formation of a surrounding debris disk that may be containing a non-vanishing dark matter fraction. In order to analyze the possible novel dark contribution, we start from a dark-matter free modellization as a benchmark. Using Monte Carlo Markov Chain (MCMC) techniques we approximately recover the gravitational waveforms, restricted to early post-merger time interval from existing simulations in the CoRe database. Later, we explore the impact of an additional dark viscous fluid under a prescribed velocity dependent force in the Lagrangian and obtain the resulting waveforms and some spectral features originating in the first few ms in the BNS post-merger. Finally we discuss our qualitative findings and its range of validity in light of the prospects of detectability in present or future experimental settings.
dark matter, effective lagrangian, gravitational waves, binary neutron star merger
@@ I Introduction
So far different particle candidates have been proposed to account for the DM. Ranging from weakly to strongly interacting species and for a mass range nowadays spanning many orders of magnitude, there are plenty of models in the literature. Just to cite a few of them, let us remind here the weakly interacting massive particle (WIMP) such as the supersymmetric neutralino, asymmetric dark matter (ADM), light bosons such as axions or axion-like particles and strangelets [6]. Although much theoretical and experimental efforts have been undertaken, its nature remains elusive, constituting an open problem. Nevertheless, there is now overwhelming evidence of its existence more than nine decades after the pioneering observations by Fritz Zwicky in 1933 [7], who showed that the velocities of galaxies in the Coma cluster greatly exceeded the expectations based solely on the sum of the individual luminous galaxy masses. Since then, wide searches involving different strategies trying to hunt these exotic particles have shown null results although the phase space is now more tightly constrained.
@@
Remarkably, the dawn of the era of multimessenger physics allows to gather information on different channels enriching the picture and more efficiently constraining the physical mechanisms governing its behaviour. In particular, gravitational waves (GWs) experiments are expected to provide a precise reconstruction of the properties of inspiraling and merging compact objects, including mass/radius/composition measurements as they would be sensitive to tiny deviations in the gravitational waveforms which may be induced by nobel features in the violent events of merging black holes (BHs) and NSs. These environmental effects could be the smoking-gun to signal the presence of DM, although this requires a careful identification of the ordinary baryon content [8].
@@
Since the LIGO first detection in 2015 of GWs emitted during the merger of a stellar-mass binary black hole (BBH) event [9], the possibility to add this new multimessenger opens a new window to the Universe. Further, GWs provide a new channel to shed light on the nature of DM, also constituting a definite confirmation that BBHs exist and merge with a local rate [10] in the range between 17.9 and $44\hskip 2.0pt\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ at a fiducial redshift $(z=0.2)$, assuming the known sources are representative of the total population. Other compact binary coalescence (CBC) search targets include collisions of BHs and NSs or binary NS mergers (BNS), where in all of them GWs are produced in a transient emission that can be detected on current terrestrial facilities. Recently, detectors including LIGO-Virgo-Kagra (LVK) have released enough data to build catalogs currently listing nearly 100 detections [11]. Specifically the Gravitational-Wave Transient Catalog 3 (GWTC-3) [10] contains signals consistent with three classes of binary mergers. They infer the BNS merger rate to be between 10 and $1700\hskip 2.0pt\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ and the NS-BH merger rate to be between 7.8 and $140\hskip 2.0pt\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, assuming a constant rate density in the comoving frame and taking the union of $90\%$ credible intervals. Besides these efforts, other future 3rd-generation planned detectors such as Cosmic Explorer (CE) [12], Einstein Telescope [13], NEMO [14] or LISA [15] will be able to further detect ultralow and ultraweak GW spectral sensitivities of the order $h\sim 10^{-25}$ $\rm Hz^{-1/2}$ produced by massive cosmic events with redshifts $z\lesssim 10$ and total mass around that expected in BNS $M\lesssim 3M_{\odot}$.
@@
In this work we will use a Lagrangian effective model inspired by Takami et al [1, 16] in order to qualitatively assess possible effects in the dynamics of the early post-merger phase in a BNS merger occurring in an environment polluted by viscous DM. Thus in our model the BNS environment is characterized by a baryonic mass fraction $f_{b}$ and DM fraction $f_{\chi}$ during the GW emission time interval. We compute the set of polarization strengths $h_{+,\times}(t)$ of the GW signal as a function of time $t$ and subsequent observables derived from them. We note that, at present, there are refined tools such as Bilby [17], a Bayesian inference library for GW astronomy, allowing to infer the source astrophysical properties from existing experimental data. This tool does not incorporate, however, particular realizations of exotic models including DM like the one we present in this work.
@@
In our setting the presence of DM in the system is due to the actual BNS merger having place in a DM polluted environment. It may be due to stripped DM from the inside of the NS or from an overdensity or halo where typically DM density can be much larger with respect to the solar neighborhood value $\rho_{\chi}\simeq 0.4$ $\rm GeV/cm^{3}$. It has been recently claimed [18] that even overdensities of a few $\sim 10^{4}$ $\rm GeV/cm^{3}$ at 1 parsec from the center of M4 would not yield definite constraints on DM from kinetically heating of NSs or white dwarfs.
@@
The detection of DM using GW along with different strategies has already been discussed, see [8]. One of the key ideas is based on the fact that it constitutes a fingerprint of the formation and evolution of first cosmic structures. Popular models include cold DM (CDM) on sub-galactic scales. However, there still remain incompatibilities between the CDM description and the observed data, as it predicts more structure on small scales than what we actually observe. Other models such as warm or fuzzy DM, as shown by simulations, lead to the suppression of small-scale structures and this, in turn, may lead to a significant reduction in the rate of BBH mergers observed, helping to further constrain among existing DM models. On the small scales, collisionless DM fluid faces major issues such as the core cusp problem, missing satellites problem or too big to fail problem, see [19]. This shows that future GW observations will provide a new probe of physics beyond the $\Lambda$CDM model [20].
@@
The possible traces that DM candidates could leave on the patterns of GW polarizations has been a particularly effervescent topic in the last decade [21, 22, 23, 24, 25]. The existence of viscous and self-interacting DM (SIDM) is another key point. At the small scale, the SIDM behaves like a collisional dark matter but on the large scale it behaves like the collisionless DM, thus fitting expectations. In [26] it is pointed out that the collisional nature of DM can lead to viscosity and using the kinetic theory, viscous coefficients of the SIDM can be obtained. Therefore the ratio of the SIDM cross section to its mass $(\sigma/m_{\chi})$, for the present observed cosmic acceleration, can be reconciled with the constraints obtained from astrophysical observations. Note that a GW will also experience the damping effect when it propagates in a fluid with non-zero viscosity so that the constraints of the luminosity distances from the observed GW events by experiments can be directly used to set constraints on self-interactions [27, 28].
@@
The structure of this contribution is as follows. In section II the Lagrangian model inspired by [1] is introduced along with the relevant set of effective parameters used to approximately reproduce the set of GW polarizations selected in our study from the Computational Relativity (CoRe) database [29] involving several equations of state (EoS). In section II.1, the different polarization modes are fit, first to a DM-free model, in order to assess the efficiency of the Monte Carlo Markov Chain (MCMC) procedure allowing for a parameter estimation and, later on, allowing for a non-vanishing viscous DM fraction. We discuss the obtained parameter values and provide a physical insight of them.
@@
In section III we calculate the power spectral density (PSD) for the full signal obtained from the selected dataset, set as a benchmark, in order to compare to that restricting to our post-merger results. This is done in order to inspect and qualitatively characterize the possible GW features and the sensitivity of current and future detectors to novel dark components, based on the relative effective signal-to-noise ratio for different DM polluted environments. Finally in section IV we conclude.
@@ II Model for the emission of Gravitational waves
The system we will scrutinize are the first few ms right after the NSs merge. Note that in this phase, strong tidal forces have already disrupted the otherwise nearly spherical (rotating) stellar configuration in asymptotically early times before the merger. In our scenario we will assume that a dark polluted environment is present due to a non-vanishing fraction of DM coexisting with the expected baryonic matter mass fraction expelled in the collision. This dark fraction origin may well come either from a liberated dark population loosely bound in superficial stellar layers or beyond baryonic radii, such as in halo configurations, see [4] or from an existing local DM overdensity if the BNS merger takes place inside a DM cluster.
As computational simulations show [30, 31, 32, 33, 34] just to cite a few, for the few ms duration the BNS early post-merger dynamics show the presence of tidal forces and in particular oscillations of the compact objects in the common gravitational potential well, characterized by the lapse function. These simulations involve complex microphysics under strong and electroweak scales in a curved space-time with dimensionality effects.
@@
Since a technical description seems a very complex task we focus on a more affordable goal of qualitatively exploring a scenario where we study how a dark fluid composed of massive $\chi$ particles may leave a fingerprint in the transient early post-merger emission of a BNS merger event. For this we will use a minimally coupled DM model. In this respect DM should be thermalized with the NS stellar matter so that it can (at least partially) follow the dynamics of the merger. This condition is fulfilled for DM models yielding constant scattering spin-independent $\chi$-nucleon cross-sections for $m_{\chi}\gtrsim 1$ GeV$/\rm c^{2}$, $\sigma_{\chi,N}\gtrsim 10^{-45}$ $\rm cm^{2}$, see [35]. Further couplings at the strong (hadronic) force limit are theoretically possible yielding a more ultradense environment [36] driving the final object to a BH collapse. However, we will restrict the DM sector we consider to yet another fraction of matter component, essentially inert with respect to changes in ordinary baryonic matter species.
II.1 Binary Neutron Star post-merger dynamical oscillations
As mentioned, we will approximately describe the early post-merger dynamical behavior as that of an spinning oscillator. A word of caution is in order regarding the scenario depicted. According to previous works [37] this model can reasonably describe BNS nearly symmetric where one of the main peaks $f_{1}$ in the spectrogram of post-merger gravitational waves is interpreted as the spin frequency. However, as claimed, this peak cannot be clearly identified for the asymmetric binary systems because the less massive NS is tidally disrupted before the merger and there are no prominent double dense cores. We assume in our calculation that the two cores survive as such during the post-merger phase i.e. the asymmetry is not large.
In that setting we will generally consider the effects of dissipation among two fluids (baryon/dark) and that from self-interacting DM. We will use an effective viscous (dissipative) force $F_{d}\sim v_{r}^{n}$ under the form of a power of radial velocity $v_{r}=\dot{r}$, see below. Calculations in other scenarios such as the global absorption signal of 21 cm line at $z\sim 17$ in the cosmic dawn era as reported in [38] propose a term to account for this gas-DM scattering terms affecting energy exchange among the two fluids $\sim\frac{dQ_{C}}{dt}$. Further, in order to have sufficient cooling of the gas at the cosmic dawn era, the DM-gas interaction is quoted to be sufficiently large, within existing constraints. We will not enter into these refinements in our model but we will consider it is non-vanishing. On general grounds in case the DM self-interaction takes the form $\sigma(v)=\sigma_{0}v^{-\alpha}$ the typical classical force in the fluid can be written as $F\sim nvp\sigma(v)=np\sigma_{0}v^{1-\alpha}$ where $n\equiv n_{\chi},p$ are density and momentum transferred in each scattering. In our setting we will restrict to integer $\alpha=[-1,1]$ but in works where BBH mergers are simulated [39] this parameter is explored, being key to facilitate the drag and collapse.
Specifically, in the case of the environment following a BNS merger, the viscosity of nuclear matter seems to be relevant at early times $\lesssim 3$ms after merger when timescales of local density oscillations and weak interactions are similar. Different neutrino microscopic approaches at the typical resolutions considered for mergers yield differences not largely affecting the early post-merger waveform, see [40].The system may undergo local density oscillations while similarly undergoing local changes in the fluid chemical composition [41].
In our scenario we will consider that the merger dynamics are affected by this non-ideal behavior with an effective viscosity from (dark/baryon) fluids including a self-interaction of the DM fraction in the form $F_{d}\simeq C_{v}\dot{r}^{n}$ with coefficients $C_{v}\equiv\alpha_{n},\beta_{n}\,(n=1,2)$ to be discussed later.
@@ II.1 Binary Neutron Star post-merger dynamical oscillations
Using MCMC we will fit the oscillation mode data as a function of time, obtained in simulations such as those in CoRe database [29] for a set of three different EoS, namely BHBlp [42], SFHo [43], and DD2 [44]. We note the dark fluid we incorporate within our Lagrangian description will nevertheless constitute a small mass fraction of the total merger mass. As discussed, the ultimate source of DM viscosity is an assimilated non-zero self-interaction and non-zero $\chi$-nucleon cross-section that although weak in strength may be, however, similar to photon-photon scattering in electromagnetic plasmas for energies or hundredths of MeV.
@@
More specifically matter will be expelled from the merging NS outer layers, less gravitationally bound, due to the tidal forces. In our setting we will consider this fraction of ejected matter forms a disk around the NS nuclei providing the scenario for Kilonova emission, being detectable with Earth based telescopes such as MAAT [45] in IR-UV bands in the neutron-rich environment formed after the merger.
@@
In this scenario we consider two merging NSs with gravitational masses $m_{1}$ and $m_{2}$ in a rotating disk with a fraction of DM, $f_{\chi}$, and an ordinary i.e. baryonic fraction, $f_{b}$, whose origin is the debris of the tidal disruption of the two compact stars. In the mechanical analog from [1], both spheres are connected by a spring characterized by constant $k$. The spheres can oscillate along the radial coordinate and rotate on the two-dimensional orbiting plane. For the sake of simplicity, we will consider an observer perpendicular to the plane where the orbital motion happens. As the damping oscillation proceeds, there will be a dynamical variation of the moment of inertia of the system and, from the conservation of angular momentum, this will be resulting in the generation of GW with a specific waveforms.
@@
Note that specifically in [1, 16] they are describing, however, a different case i.e. a four-mass oscillator where two DM and two NS matter cores oscillate jointly with two different springs with constants $k_{1},k_{2}\neq k_{1}$. Our model is different as it considers the fact that the oscillations happen in a dark environment and aims at qualitatively exploring features observed in the early post-merger signal. As we will describe, these oscillations will be affected by the presence of DM and this, in turn, determines any further changes.
Despite its simplicity, this effective model can reasonably mimic the early BNS merger signal. It provides a first-step to improve BNS early post-merger signal description, potentially enabling exploring features of GW signals in environments including a DM component.
@@
In order to evaluate the sensitivity of our model to DM in the predicted GW signal strength we will keep restricted to a minimum the number of extra parameters describing the dark environment where the BNS merger happens. Note that this effectively non-relativistic parametrization is based on the fact that there may be non-vanishing cross-sections among the two sectors and a possibly self-interacting term in the dark sector. As we will show, the non-relativistic nature of DM dynamics can be included in our minimal scenario in a meaningful way.
@@
As mentioned above, the Lagrangian we consider is able to mimic the GW polarizations of the BNS early post-merger signal. Thus the BNS merger event can be thought of in this scenario as two phases. In a first phase of the merger the outer NS layers are stripped off due to the tidal fields, carrying away the baryonic mass $m_{b}=f_{\text{ejec}}(m_{1}+m_{2})\equiv f_{\rm ejec}M_{T}$ from the progenitor stars ($m_{1},m_{2}$) rapidly forming a disk-like structure with a radius $R_{d}$, enclosing the two NS inner nuclei of masses $(1-f_{\text{ejec}})m_{1}$ and $(1-f_{\text{ejec}})m_{2}$, respectively bound by an elastic force of constant $k$. We assume the mass ratio is such that the two nuclei actually survive as such during the early post-merger time window of a few ms. We use equivalently $f_{\text{ejec}}\equiv f_{b}$. The minimum elongation happens at a characteristic radial distance, $l$, from the coordinate origin. Due to the abrupt fall of the tidal field after the ejection of the NS outer layers and the conservation of angular momentum, these nuclei keep rotating inside this pre-collapsed object. Later on, presumably a complete fusion follows inside (second phase) as soon as the rotational energy of these nuclei decays that, however, we will not study.
@@
For a system of mass ratio $q=m_{1}/m_{2}$ $(q\gtrsim 1)$, we can write the following Lagrangian as an ideal starting point